A baseball flew 19.6 m straight up into the air after being hit by a bat. How long should it take for the ball to return from its maximum height?


2.00 s


1.91 s


1.02 s


3.99 s


1
Expert's answer
2021-11-11T17:09:23-0500

"h=gt^2/2\to t=\sqrt{2h/g}=\sqrt{2\cdot 19.6/9.8}=2.00\ (s)" . Answer